OpenStudy (anonymous):

Standard form of y=-2/3x-7

OpenStudy (anonymous):

2x+3y=-21

OpenStudy (anonymous):

the std form is ax + by = c y = 2/3x -7 multiply both sides by 3 3y = 2x - 21 2x - 3y = 21

OpenStudy (anonymous):

its plus, that's a typo.